# Quantifying the Dunkelflaute: An analysis of variable renewable energy droughts in Europe

### v1

@@ Acknowledgments
Figure SI.6: Most extreme duration of single drought event of all years for each country and threshold. The latter are sorted in descending order from left to right according to the longest duration for a threshold $\tau=0.75$. In the European copperplate scenario (CP), unconstrained geographical balancing mitigates the most extreme drought duration (balancing effect).
@@
Figure SI.7: Most extreme duration of single drought events per year and threshold. Higher thresholds find long-lasting events. The year with the most extreme event duration varies across thresholds. The difference between years increases for increasing thresholds before it decreases again for very high thresholds. The ranking of years changes across thresholds. In general, combining technologies (portfolio effect) and countries (balancing effect) reduces the most extreme event duration.
SI.5 Additional illustrations of most extreme drought events
Figure SI.8 illustrates identified droughts lasting longer than one week or more, filtering out briefer events. This highlights very long-lasting below-average wind periods that may encompass multiple drought events of varying severity and solar seasonality. Combined in a portfolio, these periods are less severe (portfolio effect) and are further mitigated when assuming unconstrained geographical balancing (balancing effect). Figures SI.9SI.10, and SI.11 show identified drought patterns lasting at least one day for the years 1982/83, 2011/12, and 2013/14. Figure SI.12 additionally illustrates the most extreme drought events and optimal storage use in 1996/97 for selected regions assuming flat demand profiles, i.e., eliminating any diurnal or seasonal demand variability.
Figure SI.8: Identified drought patterns in 1996 and 1997 across all employed thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one country. To illustrate very persistent shortage situations, only droughts lasting longer than one week are displayed.
Figure SI.9: Identified drought patterns in 2011 and 2012 across all employed thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one country. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. In Spain, low PV seasonality is accompanied by wind droughts, translating into portfolio droughts arising across the turn of the year 2011/12. In contrast, due to the exceptionally high wind availability in North Europe, no portfolio droughts occur across the turn of the year in these countries.
Figure SI.10: Identified drought patterns in 2013 and 2014 across all employed thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one country. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. Brief wind droughts are infrequent in winter but longer-lasting, more severe, and more frequent in summer. Therefore, less severe portfolio droughts occur in the winter of 2013/14.
Figure SI.11: Identified drought patterns in 1982 and 1983 across all employed thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one country. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. Except for Spain, no portfolio droughts occur at the beginning of a calendar year due to the absence of wind droughts in Northern Europe.
@@
Figure SI.12: Identified most extreme drought events in 1996/97 occurring in winter (teal boxes). For countries in which the most extreme drought events occur in summer, these are additionally shown (gray boxes). For each region, portfolio drought occurrences lasting longer than one day for color-coded thresholds (upper panel) as well as exogenous flat demand profiles and optimized storage levels across three modeled interconnection scenarios (lower panel) are displayed. Note that the major storage discharging periods now perfectly coincide with the most extreme events identified by the drought mass metric in all countries, including the summer-time droughts in Sweden, Poland, and the United Kingdom.

### v2

@@ Acknowledgments
Figure SI.11: Most extreme duration of single drought event of all years for each country and across all investigated thresholds $\tau\in[0.1,1]$. The latter are sorted in descending order from left to right according to the longest duration for a threshold $\tau=0.75$. In the European copperplate scenario (CP), unconstrained geographical balancing mitigates the most extreme drought duration (balancing effect).
@@
Figure SI.12: Most extreme duration of single drought events per year and across all investigated thresholds $\tau\in[0.1,1]$. Higher thresholds find long-lasting events. The year with the most extreme event duration varies across thresholds. The difference between years increases for increasing thresholds before it decreases again for very high thresholds. The ranking of years changes across thresholds. In general, combining technologies (portfolio effect) and countries (balancing effect) reduces the most extreme event duration.
SI.6 Additional illustrations of most extreme drought events
Figure SI.13 illustrates the most extreme drought events and optimal storage use in 1996/97 for selected regions assuming flat demand profiles, i.e., eliminating any diurnal or seasonal demand variability.
@@
Figure SI.13: Identified most extreme drought events in 1996/97 occurring in winter (teal boxes). For countries in which the most extreme drought events occur in summer, these are additionally shown (gray boxes). For each region, portfolio drought occurrences lasting longer than one day for color-coded thresholds (upper panel) as well as exogenous flat demand profiles and optimized storage levels across three modeled interconnection scenarios (lower panel) are displayed. Note that the major storage discharging periods now perfectly coincide with the most extreme events identified by the drought mass metric in all countries, including the summer-time droughts in Sweden, Poland, and the United Kingdom.